OpenStudy (anonymous):

Can someone check to see if my answer is correct? Meg is starting a day-care center for the people in her community. She decided to charge $25 per day per child, out of which $5 is her expenditure per child per day. If Meg's target is to earn $2,000 over 20 weekdays in June, the minimum number of kids she has to watch per weekday that month is 5 kids per day. I say 5 because its 20 days and she need to make at least $100 a day to make $2000. But I'm not sure if it's a trick because it says weekdays and not sure if weekdays change on the calendar year to year.

OpenStudy (anonymous):

@Michele_Laino can you help please?

OpenStudy (michele_laino):

I think that every day Meg has to pay $25 per child, so, if I call with N, the number of children which Meg has to care about every day, then every day she spends 25*N

OpenStudy (michele_laino):

so we can write: \[25 \times N \leqslant 2000\]

OpenStudy (anonymous):

but it costs her $5 per day per child so she is only making $20 per child

OpenStudy (anonymous):

N would be less than or equal to 80

OpenStudy (michele_laino):

sorry, N would be 80, in only 1 day. We have to consider a time period of 20 days, so we have: \[25 \times N \times 20 \leqslant 2000\]

OpenStudy (michele_laino):

since Meg spends, in total, $25 per child, and she has only $2,000

OpenStudy (anonymous):

ok so 4 but shouldnt read 20*N*20 less than or equal to 2000 because you need to deduct $5 expenditure per child

OpenStudy (michele_laino):

yes! $25 in total

OpenStudy (anonymous):

" She decided to charge $25 per day per child, out of which $5 is her expenditure per child per day."

OpenStudy (anonymous):

that $5 isn't needed?

OpenStudy (michele_laino):

yes! I think that Meg has to charge $20 per child + $5 per expenditure of each child, so in total $25 per child

OpenStudy (anonymous):

ok so I misunderstood the question. so the answer is 4 children perday for 20 days. yes?

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (anonymous):

ok thank you!!!

OpenStudy (michele_laino):

thank you! :)

OpenStudy (anonymous):

@radar can u check this answer. The $5 expenditure leads me to believe she will only make $20 per day per child

OpenStudy (anonymous):

if there are 4 children your answer is correct

OpenStudy (anonymous):

whats that